Why is the pKa2 of H3PO4 greater than pKa1 and why is pKa3 greater than pKa2?

  1. Why is the pKa2 of H3PO4 greater than pKa1 and why is pKa3 greater than pKa2?
Answer #1

Look at basic definitions

The larger the value of Ka, the stronger the acid.

now we know that pKa = -log (Ka) ( pKa is the negative logarithm of Ka )

Therefore, the smaller value of pKa implies stronger the acid or the higher the pKa value, the weaker the acid.

Phosphoric acid H3PO4 Ka1 = 6.9×10–3
Ka2 = 6.2×10–8
Ka3 = 4.8×10–13

pKa1 is the disassociation constant for the first ionization

pKa2 is the disassociation constant for the second ionization

pKa3 is the disassociation constant for the third ionization.

pKa1 < pKa2 < pKa3

Once a proton is lost it is ok to stabilize the additional electron density through delocalization over two oxygen atoms in dihydrogen phosphate ion (H2PO4-), which is a weak acid compared to phosphoric acid. But if one more proton is lost from dihydrogen phosphate ion it develops another negative charge (-2) and creates a electrostatic repulsion between these two negative charges, moreover this charge can be delocalized on only one oxygen atom. The same concept also applicable to monohydrogen phosphate ion (HPO42-), which is also weaker acid than dihydrogen phosphate ion, leading to form more unstable phosphate ion (PO43-). The successive ionization can be shown as below

Hope this helped you!

Thank You So Much! Please Rate this answer as you wish.("Thumbs Up")